# On Viscosity Solutions of Hamilton-Jacobi Equations in the Wasserstein space and the Vanishing Viscosity Limit

### v1

@@ 4. Useful Estimates and Lemmata
The proof relies on the result contained in [CavSavSod23, Proposition  9.1] that states: In dimension $d\geq 2$, if $\mathcal{F}:\mathcal{P}_{2}(\mathbb{R}^{d})\to\mathbb{R}$ is a continuous $\Lambda$-geodesically convex function, then it is totally geodesically convex, namely given $\mu_{0},\mu_{1}\in\mathcal{P}_{2}(\mathbb{R}^{d})$ and any $\gamma\in\Gamma(\mu_{0},\mu_{1})$ we have
(4.5) $\mathcal{F}(\mu_{t})\leq(1-t)\mathcal{F}(\mu_{0})+\mathcal{F}(\mu_{1})-\frac{\Lambda}{2}t(1-t)(W_{2}^{\gamma}(\mu_{0},\mu_{1}))^{2}\quad\forall t\in[0,1],$
where $\mu_{t}=(\pi_{1}+t(\pi_{2}-\pi_{1}))_{\#}\gamma$, and $W_{2}^{\gamma}(\mu_{0},\mu_{1})^{2}=\int_{\mathbb{R}^{2d}}|y-x|^{2}d\gamma(x,y)$ is defined in (2.2).
Now, given $\nu_{0},\nu_{1}\in\mathcal{P}_{2}(\mathbb{R}^{d})$, we consider the mixture $\tau_{h,t}:=(1-h)\mu+h\nu_{t}$, defined for $h\in[0,1]$ and $t\in[0,1]$, where $\nu_{t}:=(\pi_{1}+t(\pi_{2}-\pi_{1})_{\#}\widetilde{\gamma}$ is the constant speed geodesic induced by the optimal plan $\widetilde{\gamma}\in\Gamma_{0}(\nu_{0},\nu_{1})$, and choose the plan
@@
Therefore, the total convexity property $\mathcal{F}$ yields
$\mathcal{F}(\tau_{h,t})\leq(1-t)\mathcal{F}(\tau_{h,0})+t\mathcal{F}(\tau_{h,1})-\frac{\Lambda}{2}t(1-t)(W^{\widetilde{\gamma}}_{2}(\tau_{h,0},\tau_{h,1}))^{2}.$
@@
Assertion (2) follows by choosing for all $y_{0},y_{1}\in\mathrm{supp}(\mu)$ the constant speed geodesic $\nu_{t}=\delta_{y_{t}}$, whenever $y_{t}\in\text{supp}(\mu)$, connecting $\nu_{0}=\delta_{y_{0}}$ and $\nu_{1}=\delta_{y_{1}}$: indeed in this case $\nu_{t}=\delta_{y_{t}}$ where $y_{t}=(1-t)y_{0}+ty_{1}$, and the $h\Lambda$-geodesically convexity of $\nu\mapsto\mathcal{F}((1-h)\mu+h\nu)$ restricted to Dirac masses corresponds to (4.4).
@@
We are not able to establish the previous result in dimension $d=1$, since our argument strongly relies on [CavSavSod23, Proposition 9.1], which is only available for $d>1$.
The main difficulty lies in the fact that, in general, even for $d>1$, the curve
@@
is not a $W_{2}$-geodesic, even if $t\mapsto\nu_{t}$ is one. The total convexity property avoid this issue.
@@
Suppose $d\geq 2$. Let $\mathcal{F}:\mathcal{P}_{2}(\mathbb{R}^{d})\to\mathbb{R}$ be a continous $\Lambda$-geodesically convex function. If $\mathcal{F}$ is flat differentiable at $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$ then $x\mapsto\mathcal{D}_{\mu}\mathcal{F}(\mu,x)$ is convex in $\mathrm{supp}(\mu)$.
We are now in the position to prove the announced Proposition 4.5.
@@
The argument of the proof follows the same lines of the proof of Lemma 4.3. In that context, the 1-concave function $\frac{x^{2}}{2}-\phi(x)=\mathcal{D}_{\mu}\frac{W^{2}(\mu,\nu)(x)}{2}$ is now replaced by the $\Lambda$-concave function $\mathcal{D}_{\mu}\mathcal{F}(\mu,x)$. ∎
@@
Let $U^{\varepsilon}$ be a classical subsolution of (1.2), with $a(\mu,x)=\varepsilon\mathrm{Id}_{d\times d}(x)$. If $U^{\varepsilon}$ is s.t. $\Delta_{x}\mathcal{D}_{\mu}U^{\varepsilon}(t,\mu,x)\leq C$ for some $C$ independent of $(t,\mu,x)\in[0,T]\times\mathcal{P}_{2}(\mathbb{R}^{d})\times\mathbb{R}^{d}$, then $U^{\varepsilon}-\varepsilon C(t-T)$ is a classical subsolution of (1.1). In particular, provided that $d\geq 2$, $\Delta_{x}\mathcal{D}_{\mu}U^{\varepsilon}(\mu,x)\leq C$ is satisfied by a $\Lambda$-geodesically concave functional, where $C=d\Lambda$.
@@
If $U^{\varepsilon}$ is a classical subsolution of (1.2), then $U^{\varepsilon}$ is $C^{2}$. In particular, $\Delta_{x}\mathcal{D}_{\mu}U^{\varepsilon}(\mu,x)$ is well defined in $\text{supp}(\mu)$. Moreover,
$-\partial_{t}U^{\varepsilon}+\mathcal{H}((\mathrm{Id}\times\partial_{\mu}U^{\varepsilon})_{\#}\mu)$ $=-\partial_{t}U^{\varepsilon}+\int_{\mathbb{R}^{d}}H(x,\mu,\partial_{\mu}U^{\varepsilon})d\mu$
$\underbrace{\leq}_{U^{\varepsilon}\text{classical subsolution}}\varepsilon\int_{\mathbb{R}^{d}}\Delta_{x}\mathcal{D}_{\mu}U^{\varepsilon}(\mu,x)d\mu\leq\varepsilon C.$
@@ 5. Vanishing Viscosity Limit
Let $H\colon\mathbb{R}^{d}\times\mathbb{R}^{d}\times\mathcal{P}_{2}(\mathbb{R}^{d})\to\mathbb{R}$ and $\mathcal{G}:\mathcal{P}_{2}(\mathbb{R}^{d})\to\mathbb{R}$ satisfy the assumption 2.15.
Let $U^{\varepsilon}$ and $U$ be two bounded viscosity solutions of (1.2), with $a(\mu,x)=\varepsilon\mathrm{Id}_{d\times d}(x)$, and (1.1), respectively. Assume that $U^{\varepsilon}(T,\mu)=U(T,\mu)=\mathcal{G}(\mu)$. Additionally, suppose $U^{\varepsilon}$ to be $W_{1}$-Lipschitz continuous, and $U$ to be $W_{2}$-Lipschitz continuous, Lipschitz in time, and continuous w.r.t. the weak convergence. Then there exists a constant $C\geq 0$ depending only on the data s.t.
@@
The first part of the proof is mainly inspired by the proof of [Alekos, Theorem  2.10].
Firstly, we focus on the upper bound: it will be clear through the proof that a symmetry argument, inverting the role of the two functions, allows us to concentrate on one of the two sides of convergence.
@@
Let
(5.1) $E^{+}_{\varepsilon}:=\sup_{(t,\mu)\in[0,T]\times\mathcal{P}_{2}(\mathbb{R}^{d})}U^{\varepsilon}(t,\mu)-U(t,\mu)$
be the one side error between the two solutions. We want to prove that
@@
We note that $E^{+}_{\varepsilon}$ is bounded by the boundedness assumptions on the functions. Clearly, if $E^{+}_{\varepsilon}\leq 0$ then (5.1) holds. Therefore, we assume $E^{+}_{\varepsilon}>0$.
@@
We invoke the usual doubling variables arguments by considering the penalized function
$\Phi_{\delta,\alpha}(s,t,\mu,\nu)=U^{\varepsilon}(s,\mu)-\delta(\mathcal{E}(\mu)+\pi\mathcal{M}_{2}(\mu))-U(t,\nu)-\delta\pi\mathcal{M}_{2}(\nu)-\frac{1}{2\alpha}W_{2}^{2}(\mu,\nu)-\frac{1}{2\alpha}|t-s|^{2}-\frac{2T-t-s}{4T}E^{+}_{\varepsilon}$,
@@
$\delta\mapsto M_{\delta,\alpha,\eta}:=\max_{(s,t,\mu,\nu)\in[0,T]^{2}\times(\mathcal{P}_{2}(\mathbb{R}^{d}))^{2}}\Phi_{\delta,\alpha,\eta}(s,t,\mu,\nu)$
@@
$M_{\delta,\alpha,\eta}=\Phi_{\delta,\alpha,\eta}(\overline{s},\overline{t},\overline{\mu},\overline{\nu})\leq M_{\delta/2,\alpha,\eta}-\frac{\delta}{2}(\mathcal{E}(\overline{\mu})+\pi\mathcal{M}_{2}(\overline{\mu}))-\frac{\delta\pi}{2}\mathcal{M}_{2}(\overline{\nu}).$
@@
$\frac{\delta}{2}(\mathcal{E}(\overline{\mu})+\pi\mathcal{M}_{2}(\overline{\mu}))+\frac{\delta\pi}{2}\mathcal{M}_{2}(\overline{\nu})\leq M_{\delta/2,\alpha,\eta}-M_{\delta,\alpha,\eta}\xrightarrow{\delta\to 0}0$
for fixed $\alpha,\eta>0$. In particular,
@@
Now, rearranging the terms in the inequality $\Phi_{\delta,\alpha,\eta}(\overline{s},\overline{t},\overline{\mu},\overline{\mu})\leq\Phi_{\delta,\alpha,\eta}(\overline{s},\overline{t},\overline{\mu},\overline{\nu})$, leads to the inequality
@@
By the $W_{2}$-Lipschitz regularity of $U$, we recover the following bound
@@
which implies, thanks to the Lipschitz continuity in time of $U$ and the boundedness of $E_{\varepsilon}^{+}$,
@@
By adapting Lemma 4.2 to functions dependent on time,
@@
where $\overline{\Sigma}=(\pi_{1},\frac{\pi_{1}-\pi_{2}}{\alpha}-2\pi\delta\pi_{2})_{\#}\overline{\sigma}$, with $\overline{\sigma}\in\Gamma_{0}(\overline{\nu},\overline{\mu})$ s.t. $(\pi_{2},\pi_{1})_{\#}\overline{\sigma}=(\mathrm{Id}\times T_{\overline{\mu}}^{\overline{\nu}})_{\#}\overline{\mu}$. Moreover, we have $\mathcal{I}(\overline{\mu})<\infty$. Therefore, we can apply the Definitions 2.18 and 2.16 to obtain the following inequalities
@@
Using Cauchy-Schwarz inequality and (5.2), (5.3) we get
@@
The third integral $I_{3}$ can be bound thanks to the fact that $\overline{\mu}\in W_{\mathrm{loc}}^{1,\infty}$and $\overline{\mu}^{-1}\in L_{\mathrm{loc}}^{\infty}$. Indeed, in this case, as in the proof in [DS, Theorem  2.14], we can perform an integration by parts and prove
@@
$E_{\varepsilon}^{-}:=\sup_{(t,\mu)\in[0,T]\times\mathcal{P}_{2}(\mathbb{R}^{d})}U(t,\mu)-U^{\varepsilon}(t,\mu)\leq C\sqrt{\varepsilon}.$
@@
The boundedness assumption on the functions can be relaxed to allow for quadratic growth, as already observed in [DS]. Indeed, in the proof it suffices to introduce an additional moment penalization to compensate for the growth of the difference $U^{\varepsilon}-U$.
@@
As in the finite-dimensional setting, the rate of convergence follows from the regularity assumptions on the Hamiltonian: modifying its local modulus of continuity directly affects the convergence speed. The argument can be adapted to this framework without additional difficulty. Finally, we can also handle the case of terminal costs depending on the viscosity parameter $\varepsilon$ thanks to the stability property established in [DS, Theorem  4.1].
@@
The above method yields an optimal convergence rate. For a Hamiltonian satisfying 2.15, this rate is the best one can expect without assuming convexity of $H$, since the finite-dimensional setting is embedded in this framework. It would be interesting to investigate whether this bound can be improved in the convex case, as independently shown in [cirant2025convergenceratesvanishingviscosity, chaintron2025optimalrateconvergencevanishing].
@@
Under the assumption of Theorem 5.1, if $U^{\varepsilon}$ is a classical solution of (1.2) s.t. $\Delta\mathcal{D}_{\mu}U^{\varepsilon}\leq C$ where $C:[0,T]\to[0,+\infty)$ Then
$U^{\varepsilon}(t,\mu)-U(t,\mu)\leq C(T-t)\varepsilon.$
@@
It is sufficient to observe that we can apply Lemma 4.11 and then the comparison principle in [BSOT, Theorem  2.11]. ∎

### v2

@@ 4. Useful Estimates and Lemmata
The proof relies on the result contained in [15, Proposition  9.1] that states: In dimension $d\geq 2$, if $\mathcal{F}:\mathcal{P}_{2}(\mathbb{R}^{d})\to\mathbb{R}$ is a continuous $\Lambda$-geodesically convex function, then it is totally geodesically convex, namely given $\mu_{0},\mu_{1}\in\mathcal{P}_{2}(\mathbb{R}^{d})$ and any $\gamma\in\Gamma(\mu_{0},\mu_{1})$ we have
$\mathcal{F}(\mu_{t})\leq(1-t)\mathcal{F}(\mu_{0})+\mathcal{F}(\mu_{1})-\frac{\Lambda}{2}t(1-t)(W_{2}^{\gamma}(\mu_{0},\mu_{1}))^{2}\quad\forall t\in[0,1],$
where $\mu_{t}=(\pi_{1}+t(\pi_{2}-\pi_{1}))_{\#}\gamma$, and $(W_{2}^{\gamma}(\mu_{0},\mu_{1}))^{2}:=\int_{\mathbb{R}^{2d}}|y-x|^{2}d\gamma(x,y)$ is defined in (2.2).
Now, let $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$, given $\nu_{0},\nu_{1}\in\mathcal{P}_{2}(\mathbb{R}^{d})$, we consider the mixture $\tau_{h,t}:=(1-h)\mu+h\nu_{t}$, defined for $h\in[0,1]$ and $t\in[0,1]$, where $\nu_{t}:=(\pi_{1}+t(\pi_{2}-\pi_{1}))_{\#}\widetilde{\gamma}$ is the $W_{2}$-constant speed geodesic induced by the optimal plan $\widetilde{\gamma}\in\Gamma_{0}(\nu_{0},\nu_{1})$, and choose the plan
@@
Therefore, the total convexity property of $\mathcal{F}$ yields
$\mathcal{F}(\tau_{h,t})\leq(1-t)\mathcal{F}(\tau_{h,0})+t\mathcal{F}(\tau_{h,1})-\frac{\Lambda}{2}t(1-t)(W^{\widetilde{\gamma}}_{2}(\tau_{h,0},\tau_{h,1}))^{2},$
@@
Assertion (2) follows by choosing for all $y_{0},y_{1}\in\mathbb{R}^{d}$ the $W_{2}$-constant speed geodesic $\nu_{t}=\delta_{y_{t}}$, where $y_{t}=(1-t)y_{0}+ty_{1}$, connecting $\nu_{0}=\delta_{y_{0}}$ and $\nu_{1}=\delta_{y_{1}}$: indeed in this case $\nu_{t}=\delta_{y_{t}}$, and the $h\Lambda$-geodesically convexity of $\nu\mapsto\mathcal{F}((1-h)\mu+h\nu)$ restricted to Dirac masses corresponds to (4.3).
@@
Note that the main difficulty in the proof lies in the fact that, in general, the curve
@@
is not necessarily a $W_{2}$-constant speed geodesic, even if $t\mapsto\nu_{t}$ is one. The total convexity property, proved when $d\geq 2$ thanks to [15, Proposition 9.1], avoid this issue. Being our argument strongly based on that Proposition, we couldn’t remove the constraint $d\geq 2$.
As a byproduct of the previous lemmas, we obtain the following corollary.
@@
Suppose $d\geq 2$. Let $\mathcal{F}:\mathcal{P}_{2}(\mathbb{R}^{d})\to\mathbb{R}$ be a continuous $\Lambda$-geodesically convex function. If $\mathcal{F}$ is flat differentiable at $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$ then $\mathbb{R}^{d}\ni x\mapsto\mathcal{D}_{\mu}\mathcal{F}(\mu,x)$ is $\Lambda$-convex
We are now ready to prove the announced Proposition 4.7.
@@
The argument of the proof follows the same lines of the proof of Lemma 4.4. In that context, the 1-concave function $\frac{x^{2}}{2}-\phi(x)=\mathcal{D}_{\mu}\frac{W^{2}(\cdot,\nu)}{2}(\mu,x)$ is now replaced by the $\Lambda$-concave function $x\mapsto\mathcal{D}_{\mu}\mathcal{F}(\mu,x)$. ∎
@@
Let $U^{\varepsilon}$ be a classical subsolution of (1.2), with $a(\mu,x)=\varepsilon\mathrm{Id}_{d\times d}$. If $U^{\varepsilon}$ is s.t. $\Delta_{x}\mathcal{D}_{\mu}U^{\varepsilon}(t,\mu,x)\leq C$ for some $C$ independent of $(t,\mu,x)\in[0,T]\times\mathcal{P}_{2}(\mathbb{R}^{d})\times\mathbb{R}^{d}$, then $U^{\varepsilon}-\varepsilon C(T-t)$ is a classical subsolution of (1.1). In particular, the inequality $\Delta_{x}\mathcal{D}_{\mu}U^{\varepsilon}(t,\mu,x)\leq C$ is satisfied with $C=d\Lambda$, provided that $d\geq 2$ and $U^{\varepsilon}$ is a $\Lambda$-geodesically concave functional.
@@
If $U^{\varepsilon}$ is a classical subsolution of (1.2), then $U^{\varepsilon}$ is partially $C^{2}$. In particular, $\Delta_{x}\mathcal{D}_{\mu}U^{\varepsilon}(\mu,x)$ is well defined in $\text{supp}(\mu)$. Moreover,
$-\partial_{t}U^{\varepsilon}+\mathcal{H}((\mathrm{Id}\times\partial_{\mu}U^{\varepsilon})_{\#}\mu)$ $=-\partial_{t}U^{\varepsilon}+\int_{\mathbb{R}^{d}}H(x,\partial_{\mu}U^{\varepsilon},\mu)d\mu(x)$
$\underbrace{\leq}_{U^{\varepsilon}\text{classical subsolution}}\varepsilon\int_{\mathbb{R}^{d}}\Delta_{x}\mathcal{D}_{\mu}U^{\varepsilon}(t,\mu,x)d\mu(x)\leq\varepsilon C.$
@@ 5. Vanishing Viscosity Limit
Suppose Assumptions 2.16 hold. Let $U^{\varepsilon}$ and $U$ be two bounded viscosity solutions of (1.2) and (1.1), respectively, s.t. that $U^{\varepsilon}(T,\mu)=U(T,\mu)=\mathcal{G}(\mu)$. Additionally, suppose $U^{\varepsilon}$ to be uniformly in time $W_{1}$-Lipschitz continuous for each $\varepsilon$, and $U$ to be uniformly in time $W_{2}$-Lipschitz continuous. Moreover, suppose $U$ to be jointly continuous with respect to the product topology of $0,T]\times\mathcal{P}_{2}(\mathbb{R}^{d})$, where $\mathcal{P}_{2}(\mathbb{R}^{d})$ is endowed with the weak convergence as in Definition 2.17. Then, there exists a constant $C\geq 0$, depending only on the data, such that
@@
The first part of the proof is mainly inspired by the proof of [17, Theorem  2.10].
Firstly, we focus on the upper bound
(5.1) $E^{+}_{\varepsilon}:=\sup_{(t,\mu)\in[0,T]\times\mathcal{P}_{2}(\mathbb{R}^{d})}U^{\varepsilon}(t,\mu)-U(t,\mu).$
It will be clear through the proof that a symmetry argument, inverting the role of the two functions, will give also the bound for
$E_{\varepsilon}^{-}:=\sup_{(t,\mu)\in[0,T]\times\mathcal{P}_{2}(\mathbb{R}^{d})}U(t,\mu)-U^{\varepsilon}(t,\mu).$
@@
We want to prove that
@@
We note that $E^{+}_{\varepsilon}$ is bounded by the boundedness assumptions on the two functions $^{\varepsilon}, UU$. Clearly, if $E^{+}_{\varepsilon}\leq 0$ then (5.1) holds. Therefore, we assume $E^{+}_{\varepsilon}>0$.
@@
We invoke the usual doubling variables argument by considering the penalized function
$\Phi_{\delta,\alpha}(s,t,\mu,\nu):=$ $\,U^{\varepsilon}(s,\mu)-\delta(\mathcal{E}(\mu)+\pi\mathcal{M}_{2}(\mu))-U(t,\nu)-\delta\pi\mathcal{M}_{2}(\nu)-\frac{1}{2\alpha}W_{2}^{2}(\mu,\nu)$
$\!-\frac{1}{2\alpha}|t-s|^{2}-\frac{2T-t-s}{4T}E^{+}_{\varepsilon}$,
@@
By adapting Lemma (4.3) to functions dependent on time, the function $\Phi_{\delta,\alpha}$ attains its maximum. We denote by $(\overline{s},\overline{t},\overline{\mu},\overline{\nu})\in[0,T]^{2}\times\mathcal{P}_{2}(\mathbb{R}^{d})^{2}$ a point where this maximum is achieved.
@@
$\delta\mapsto M_{\delta,\alpha}:=\max_{(s,t,\mu,\nu)\in[0,T]^{2}\times(\mathcal{P}_{2}(\mathbb{R}^{d}))^{2}}\Phi_{\delta,\alpha}(s,t,\mu,\nu)$
@@
$M_{\delta,\alpha}=\Phi_{\delta,\alpha}(\overline{s},\overline{t},\overline{\mu},\overline{\nu})\leq M_{\delta/2,\alpha}-\frac{\delta}{2}(\mathcal{E}(\overline{\mu})+\pi\mathcal{M}_{2}(\overline{\mu}))-\frac{\delta\pi}{2}\mathcal{M}_{2}(\overline{\nu}).$
@@
$\frac{\delta}{2}(\mathcal{E}(\overline{\mu})+\pi\mathcal{M}_{2}(\overline{\mu}))+\frac{\delta\pi}{2}\mathcal{M}_{2}(\overline{\nu})\leq M_{\delta/2,\alpha}-M_{\delta,\alpha}\xrightarrow{\delta\to 0}0$
for fixed $\alpha>0$. In particular,
@@
Now, rearranging the terms in the inequality $\Phi_{\delta,\alpha}(\overline{s},\overline{t},\overline{\mu},\overline{\mu})\leq\Phi_{\delta,\alpha}(\overline{s},\overline{t},\overline{\mu},\overline{\nu})$, leads to the inequality
@@
By the uniform in time $W_{2}$-Lipschitz regularity of $U$, we recover the following bound
@@
which implies, thanks to the uniform Lipschitz continuity in time of $U$, ensured by Lemma 4.2, and the boundedness of $E_{\varepsilon}^{+}$,
@@
By adapting Lemma 4.3 to functions dependent on time, it follows
@@
where $\overline{\Sigma}=(\pi_{1},\frac{\pi_{1}-\pi_{2}}{\alpha}-2\pi\delta\pi_{2})_{\#}\overline{\sigma}$, with $\overline{\sigma}\in\Gamma_{0}(\overline{\nu},\overline{\mu})$ s.t. $(\pi_{2},\pi_{1})_{\#}\overline{\sigma}=(\mathrm{Id}\times T_{\overline{\mu}}^{\overline{\nu}})_{\#}\overline{\mu}$. Moreover, we have $\mathcal{I}(\overline{\mu})<\infty$. Therefore, we can apply the Definitions 2.21 and 2.19 to obtain the following inequalities
@@
Using Cauchy Schwarz inequality and (5.2), (5.3) we get
@@
The third integral $I_{3}$ can be bounded thanks to the fact that $\overline{\mu}\in W_{\mathrm{loc}}^{1,\infty}$and $\overline{\mu}^{-1}\in L_{\mathrm{loc}}^{\infty}$. Indeed, in this case, as in the proof in [30, Theorem  2.14], we can perform an integration by parts and prove
@@
$E_{\varepsilon}^{-}\leq C\sqrt{\varepsilon}.$
@@
The boundedness assumption on $U$ and $U^{\varepsilon}$ can be relaxed to a more general growth condition, as already observed in [30]. We also observe that the additional continuity of $U$ w.r.t. the weak convergence can be dropped in the case the equation is considered in $\mathcal{P}(\mathbb{T}^{d})$.
@@
As in the finite-dimensional setting, the rate of convergence follows from the regularity assumptions on the Hamiltonian: modifying its local modulus of continuity directly affects the convergence speed. The argument can be adapted to this framework without additional difficulty. Finally, we can also handle the case of terminal costs depending on the viscosity parameter $\varepsilon$ thanks to the stability property established in [30, Theorem  4.1].
@@
The above method yields an optimal convergence rate. For a Hamiltonian satisfying Assumption 2.16, this rate is the best one can expect without assuming convexity of $H$, since the finite-dimensional setting is embedded in this framework. It would be interesting to investigate whether this bound can be improved in the convex case, as independently shown in [19, 18].
@@
Under the assumption of Theorem 5.1, if $U^{\varepsilon}$ is a classical subsolution of (1.2) s.t. $\Delta_{x}\mathcal{D}_{\mu}U^{\varepsilon}(t,\mu,x)\leq C$ for some $C$ independent of $(t,\mu,x)\in[0,T]\times\mathcal{P}_{2}(\mathbb{R}^{d})\times\mathbb{R}^{d}$. Then
$U^{\varepsilon}(t,\mu)-U(t,\mu)\leq C(T-t)\varepsilon,\quad\forall(t,\mu)\in[0,T]\times\mathcal{P}_{2}(\mathbb{R}^{d}).$
@@
It is sufficient to observe that we can apply Lemma 4.12 and then the comparison principle in [9, Theorem  2.11]. ∎